OpenStudy (anonymous):

Given the following function: f(x) = { |3+x|/3+x if x < -3 x + K if -3 4 Determine the lim x-> 4+ f(x) I am not sure how to approach this problem.

myininaya (myininaya):

\[\lim_{x \rightarrow 4^+}f(x)=\lim_{x \rightarrow 4^+}(\frac{\sqrt{x}-2}{x-4})\]

myininaya (myininaya):

you can factor the bottom

myininaya (myininaya):

\[x-4=(\sqrt{x}-2)(\sqrt{x}+2)\]

myininaya (myininaya):

any questions? I looked at the part of function where x>4 because we wanted to look to the right of 4

OpenStudy (anonymous):

so we would get 1/4 as a result?

myininaya (myininaya):

yes \[\lim_{x \rightarrow 4}\frac{\sqrt{x}-2}{(\sqrt{x}-2)(\sqrt{x}+2)}=\lim_{x \rightarrow 4} \frac{1}{ \sqrt{x}+2}\]

myininaya (myininaya):

1/(2+2) is 1/4

OpenStudy (anonymous):

wow! thanks so much, this makes way more sense now.

myininaya (myininaya):

great

OpenStudy (anonymous):

Great profile picture! @Zarkon

OpenStudy (anonymous):

lim f(x)

OpenStudy (anonymous):

same answer as my prof wrote on the exam.. just didn't understand how he got it!

OpenStudy (zarkon):

@No-data ty